# Generalized cryptographic semi-regular sequences: A variant of Fröberg conjecture and a simple complexity estimation for Gröbner basis computation

### v1

@@ 2.3 Genericness and a variant of Fröberg’s conjecture
Another definition of genericness is the following (cf. [20, page 119]): For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, a property $\mathcal{P}$ for $\bm{F}=(f_{1},\ldots,f_{m})\in V$ is said to be generic over $K$ if it holds for any $\bm{F}=(f_{1},\ldots,f_{m})\in V$ such that the coefficients of $f_{1},\ldots,f_{m}$ are algebraically independent over the prime field $L$ of $K$, that is, $L=\mathbb{Q}$ or $\mathbb{F}_{p}$ for some prime number $p$. It can be checked that the two definitions of genericness are equivalent.
@@
Assume that $K$ is an infinite field. For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, the property that $\bm{F}=(f_{1},\ldots,f_{m})\in V$ satisfies (2.2.2) (i.e., $\bm{F}$ is cryptographic semi-regular) is generic. In other words, a generic sequence $\bm{F}=(f_{1},\ldots,f_{m})\in V$ is cryptographic semi-regular.
See Pardue’s paper [34] for some conjectures equivalent to Fröberg’s one (Conjecture 2.3.3). He also proved that Moreno-Socías conjecture [33] is stronger than Conjecture 2.3.3, see [34, Theorem 2] for a proof.
Here, as a variant of Conjecture 2.3.3, we raise the following conjecture (cf. [29, Conjecture 4.3.4] as its strict version):
Conjecture 2.3.4.
Assume that $K$ is an infinite field. For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, the property that $\bm{F}=(f_{1},\ldots,f_{m})\in V$ is generalized cryptographic semi-regular is generic.
We shall prove Conjectures 2.3.3 and 2.3.4 are equivalent to each other. For this, we first prove the following lemma:
@@
For $m\geq n$, the property that $R/\langle f_{1},\ldots,f_{m}\rangle$ is Artinian is generic.
@@
The case $m=n$ is well-known to hold. Assume $m>n$, fix $n$, $m$, and $(d_{1},\ldots,d_{m})$, and put
$\Lambda^{\rm Art}:=\{\bm{F}=(f_{1},\ldots,f_{m})\in V:\mbox{$R/\langle f_{1},\ldots,f_{m}\rangle$ is Artinian}\}.$
Consider a natural projection
$V\to V_{n}:=\mathbb{A}^{N_{1}}(K)\times\dots\times\mathbb{A}^{N_{n}}(K)\ ;\ (f_{1},\ldots,f_{n},f_{n+1},\ldots,f_{m})\mapsto(f_{1},\ldots,f_{n}).$
Then we have $\pi^{-1}(\Lambda_{n}^{\rm Art})\subset\Lambda^{\rm Art}$, where
$\Lambda_{n}^{\rm Art}:=\{\bm{F}=(f_{1},\ldots,f_{n})\in V:\mbox{$R/\langle f_{1},\ldots,f_{n}\rangle$ is Artinian}\}$
is Zariski-open in $V_{n}$. Since $\pi$ is continuous, $\pi^{-1}(\Lambda_{n}^{\rm Art})$ is Zariski-open in $V$, whence the assertion holds. ∎
Proposition 2.3.6.
Assume that $K$ is an infinite field. For $m\geq n$, our conjecture (Conjecture 2.3.4) is equivalent to Fröberg’s conjecture (Conjecture 2.3.3).
Proof.
Fixed $n$, $m$, and $(d_{1},\ldots,d_{m})$, we set
[…]
so that $\Lambda^{\rm csr}=\Lambda^{\rm gcsr}\cap\Lambda^{\rm Art}$.
Assume that $(f_{1},\ldots,f_{m})\in V$ is generically generalized cryptograhic semi-regular, i.e., there exists a non-empty Zariski-open set $U_{1}\subset V$ such that $U_{1}\subset\Lambda^{\rm gcsr}$. Recall from Lemma 2.3.5 that $(f_{1},\ldots,f_{m})\in V$ generically has the coordinate ring $R/\langle f_{1},\ldots,f_{m}\rangle$ that is Artinian, whence there exists a non-empty Zariski-open set $U_{2}\subset V$ such that $U_{2}\subset\Lambda^{\rm Art}$. Since both $U_{1}$ and $U_{2}$ are dense, we have $\emptyset\neq U_{1}\cap U_{2}\subset\Lambda^{\rm gcsr}\cap\Lambda^{\rm Art}=\Lambda^{\rm csr}$. Hence, $(f_{1},\ldots,f_{m})\in V$ is cryptographic semi-regular on the non-empty Zariski open set $U_{1}\cap U_{2}$ of $V$, as desired.
The converse is obvious by $\Lambda^{\rm csr}\subset\Lambda^{\rm gcsr}$. ∎
Despite of the equivalence between two conjectures, those cannot guarantee the genericness of $\Lambda^{\rm gscr}\cap(V\smallsetminus\Lambda^{\rm Art})$ with respect to the induced topology of $V\smallsetminus\Lambda^{\rm Art}$. Once the genericness holds for the case, we have the following.
Remark 2.3.7.
If the subset $\Lambda^{\rm gcsr}\cap(V\smallsetminus\Lambda^{\rm Art})$ contains a non-empty Zariski-open set in the subspace $V\smallsetminus\Lambda^{\rm Art}$ equipped with the induced topology, then ‘almost’ all sequences $(f_{1},\ldots,f_{m})$ such that $R/\langle f_{1},\ldots,f_{m}\rangle$ has Krull dimension $\geq 1$ are expected to be generalized cryptographic semi-regular.
As the first discussion about the genericness of $\Lambda^{\rm gcsr}\cap(V\smallsetminus\Lambda^{\rm Art})$, we consider the following sub case: Let $Z$ be the non-empty Zariski closed set in $V=V(n,m,d_{1},\ldots,d_{m})$ defined by the condition that the $x_{n}^{d_{i}}$-coefficient of $f_{i}$ is zero for any $i$ with $1\leq i\leq m$. Then, for any element $\bm{F}=(f_{1},\ldots,f_{m})\in Z$, the homogeneous ideal $\langle f_{1},\ldots,f_{m}\rangle_{R}$ has a projective root $(0:\cdots:0:1)$, whence $R/\langle f_{1},\ldots,f_{m}\rangle$ is not Artinian. Namely, we have $Z\subset V\smallsetminus\Lambda^{\rm Art}$.
Lemma 2.3.8.
Let $\bm{G}=(g_{1},\ldots,g_{m})\in Z$, and assume that the other coefficients of $g_{1},\ldots,g_{m}$ are algebraically independent over the prime field of $K$. Then, for any $\bm{F}=(f_{1},\ldots,f_{m})\in Z$, we have the coefficient-wise inequality
${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}(z)\geq{\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}(z).$
Therefore, if there exists a generalized cryptographic semi-regular sequence $\bm{F}=(f_{1},\ldots,f_{m})$ in $Z$ with $\dim_{K}(R/\langle f_{1},\ldots,f_{m}\rangle)_{D}=1$ for $D=\tilde{d}_{\rm reg}(\langle f_{1},\ldots,f_{m}\rangle_{R})$, then $\bm{G}=(g_{1},\ldots,g_{m})$ is generalized cryptographic semi-regular, i.e.,
${\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}(z)\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D^{\prime}}}$
for $D^{\prime}=\widetilde{d}_{\rm reg}(\langle g_{1},\ldots,g_{m}\rangle_{R})$, and $\dim_{K}(R/\langle g_{1},\ldots,g_{m}\rangle)_{D^{\prime}}=1$.
Proof.
The first half can be proved in a way similar to the proof of [20, Lemma 1]. To prove the second one, we also use Lemma 2.2.5 (cf. [20]) for a sequence $\bm{H}=(h_{1},\ldots,h_{m})$ of generic forms with the same parameters $n,m,d_{1},\ldots,d_{m}$. Then, we have ${\rm HS}_{R/\langle\bm{G}\rangle}\geq{\rm HS}_{R/\langle\bm{H}\rangle}$ by [20, Lemma 1] and thus,
${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}\geq{\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}\geq{\rm HS}_{R/\langle h_{1},\ldots,h_{m}\rangle}\geq^{\rm lex}\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}\right],$ (2.3.1)
where $\geq^{\rm lex}$ denotes the lexicographical inequality. As $\bm{F}$ is a generalized cryptographic semi-regular sequence, we also have
${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D}}$ (2.3.2)
for $D=\tilde{d}_{\rm reg}(\langle f_{1},\ldots,f_{m}\rangle_{R})$. Combining (2.3.1) and (2.3.2), we have
${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}\equiv{\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D}}.$
Meanwhile, it follows from the condition $\dim_{K}(R/\langle f_{1},\ldots,f_{m}\rangle)_{D}=1$ that we obtain $\dim_{K}(R/\langle g_{1},\ldots,g_{m}\rangle)_{D}=1$ and $\tilde{d}_{\rm reg}(\langle g_{1},\ldots,g_{m}\rangle_{R})\leq D$. Hence we have
${\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D^{\prime}}}$
for $D^{\prime}=\tilde{d}_{\rm reg}(\langle g_{1},\ldots,g_{m}\rangle_{R})$. ∎
From Lemma 2.3.8 and Remark 2.3.2, if there exists a generalized cryptographic semi-regular sequence $\bm{F}=(f_{1},\ldots,f_{m})$ with $\dim_{K}(R/\langle f_{1},\ldots,f_{m}\rangle)_{D}=1$ for $D=\tilde{d}_{\rm reg}(\langle f_{1},\ldots,f_{m}\rangle_{R})$, it follows that $Z\cap\Lambda^{\rm gcsr}$ contains a non-empty Zariski open set in $Z$ equipped with the induced topology, which implies that [29, Conjecture 4.3.4] is true for the parameter $(n,m,d_{1},\ldots,d_{m})$.
3 Proofs of main results
Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$ (in any characteristic), and let $I$ be a proper homogeneous ideal of $R$. Let $F=\{f_{1},\ldots,f_{m}\}$ be a subset of $R$ generating $I$, and put $d_{j}:=\deg(f_{j})$ (total degree) for $1\leq j\leq m$.
For a linear transformation $\sigma$ of variables $x_{1},\ldots,x_{n}$ over $K$ represented by a $n\times n$ matrix $P$ over $K$, we denote by $h^{\sigma}:=h(\bm{x}\cdot P)$ the image of $h\in R$ by $\sigma$, where $\bm{x}:=(x_{1},\ldots,x_{n})$.
3.1 Proofs of Theorem 1 (1) and (2)
Let $\prec$ be the graded reverse lexicographic ordering on $R$ with $x_{n}\prec\cdots\prec x_{1}$. We denote by $N$ the number of its projective zeros over $\overline{K}$ of the homogeneous ideal $I\subset R$. Throughout this subsection, assume that $\mathrm{Krull.dim}(R/I)\leq 1$, i.e., $N<\infty$.
Lemma 3.1.1.
If the cardinality of $K$ is greater than $N$, there exist a non-negative integer $d$ and a linear form $\ell\in R$ with $\ell\notin I$ such that the $K$-linear map $(R/I)_{d-1}\to(R/I)_{d}$ defined by the multiplication by $\ell$ is surjective.
Proof.
The assertion is proved in the proof of (i) $\Rightarrow$ (iv) in [11, Theorem 3.3.4] when $K$ is an infinite field: In fact, it suffices for the proof that one can take a linear form $\ell\in R$ that does not vanish at any projective zero over $\overline{K}$ of $I$. Even in the case where $K$ is an finite field $\mathbb{F}_{q}$, if $q$ is greater than the number $N$ of projective zeros over $\overline{K}$ of $I$, we prove that such a linear form always exists.
For each projective zero $\bm{a}=(a_{1}:\cdots:a_{n})\in\mathbb{P}^{n-1}(\overline{K})$ of $I$, we denote by $S_{\bm{a}}$ the set of all linear forms that vanish at $\bm{a}$, say
$S_{\bm{a}}=\{y_{1}x_{1}+\cdots+y_{n}x_{n}\mid y_{i}\in K\mbox{ and }y_{1}a_{1}+\cdots y_{n}a_{n}=0\},$
$S_{\bm{a}}^{\prime}=\{(y_{1},\ldots,y_{n})\in K^{n}\mid a_{1}y_{1}+\cdots a_{n}y_{n}=0\}.$
The dimension of $S_{\bm{a}}^{\prime}$ as a $K$-linear space is at most $n-1$. Indeed, for an arbitrary non-zero entry $a_{i}$ of $\bm{a}$, the vector with $1$ in the $i$-th entry and 00’s elsewhere does not belong to $S_{\bm{a}}^{\prime}$, so that $S_{\bm{a}}^{\prime}$ is a proper subspace of $K^{n}$.
Here, we have
#$\#\left(\bigcup_{\bm{a}\in V_{\overline{K}}(I)}S_{\bm{a}}\right)\leq\sum_{{\bm{a}\in V_{\overline{K}}(I)}}\#S_{\bm{a}}\leq Nq^{n-1}<q^{n},$
whence a desired linear form over $K$ exists. ∎
Proof of Theorem 1 (1):
Note that, by a choice of $\ell$, the Krull dimension of $R/\langle I,\ell\rangle$ is zero. For $d^{\prime}\geq 1$, the $-\ell$ map $(R/I)_{d^{\prime}-1}\to(R/I)_{d^{\prime}}$ is surjective if and only if $(R/I)_{d^{\prime}}=0$, which follows from the following exact sequence:
${(R/I)_{d^{\prime}-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\times\ell}$${(R/I)_{d^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(R/\langle I,\ell\rangle)_{d^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
When $d^{\prime}=d$, we have $(R/\langle I,\ell\rangle)_{d}=0$, by the surjectivity of the multiplication map $(R/I)_{d^{\prime}-1}\to(R/I)_{d^{\prime}}$ by $\ell$. Thus, it follows that $(R/\langle I,\ell\rangle)_{d^{\prime}}=0$ for any $d^{\prime}\geq d$, from which the $-\ell$ map $(R/I)_{d^{\prime}-1}\to(R/I)_{d^{\prime}}$ is surjective for any $d^{\prime}$ with $d^{\prime}\geq d$. Note that the minimal $d$ for which the map is surjective is equal to $d_{\rm reg}(\langle I,\ell\rangle_{R})$.
We first note that $\widetilde{d}_{\rm reg}(I)=\widetilde{d}_{\rm reg}(I^{\tau})$ and ${d}_{\rm reg}(\langle I,\ell\rangle_{R^{\prime}})={d}_{\rm reg}(\langle I^{\tau},\ell^{\tau}\rangle_{R^{\prime}})$ for any invertible linear transformation $\tau$ of variables. There exists a $k\in\{1,\ldots,n\}$ such that $x_{k}$-coefficient of $\ell$ is not zero. We may also assume that the $x_{k}$-coefficient is $1$, without loss of generality. Let $\sigma_{1}$ be the linear transformation of variables $x_{1},\ldots,x_{n}$ defined by a permutation sending $x_{k}$ to $x_{n}$, and write $\ell^{\sigma_{1}}=a_{0}x_{0}+a_{1}x_{1}+\cdots+a_{n}x_{n}$ for $a_{i}\in K$ with $a_{n}=1$. Then, let $\sigma_{2}$ be a linear transformation of variables $x_{1},\ldots,x_{n}$ represented by the matrix
$P:=\begin{pmatrix}1&&&-a_{1}\\ &\ddots&&\vdots\\ &&1&-a_{n-1}\\ &&&1\\ \end{pmatrix},$
where blank entries are 00’s. Namely, for $h\in R$, we define
$h^{\sigma_{2}}:=h(\bm{x}\cdot P)=h(x_{1},\ldots,x_{n-1},x_{n}-a_{1}x_{1}-\cdots-a_{n-1}x_{n-1})$
with $\bm{x}:=(x_{1},\ldots,x_{n})$. Then, the invertible transformation $\sigma:=\sigma_{2}\circ\sigma_{1}$ of variables $x_{1},\ldots,x_{n}$ sends $\ell$ to $x_{n}$, and we have the equalities $D^{\prime}:=\widetilde{d}_{\rm reg}(I)=\widetilde{d}_{\rm reg}(I^{\sigma})$ and $D:=d_{\rm reg}(\langle I,\ell\rangle_{R})=d_{\rm reg}(\langle I^{\sigma},x_{n}\rangle_{R})$. The assertion follows from Theorem 2.1.9.∎
In the following, we shall prove Theorem 1 (2), for which we first prove the following easy lemma in the case where $m=n-1$:
Lemma 3.1.2.
With notation as above, if $m=n-1$ and if the sequence $(f_{1},\ldots,f_{m},\ell)$ is cryptographic semi-regular (equivalently, it is regular since $m+1=n$), then we have $\widetilde{d}_{\rm reg}(I)=d_{\rm reg}(\langle I,\ell\rangle_{R})-1$, and hence
$\mathrm{max.GB.deg}_{\prec}(I^{\sigma})\leq\mathrm{max}\{d_{\rm reg}(\langle I,\ell\rangle_{R}),\widetilde{d}_{\rm reg}(I)\}=d_{\rm reg}(\langle I,\ell\rangle_{R})=\sum_{j=1}^{n-1}(d_{j}-1)+1.$
Moreover, $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular (equivalently, it is regular since $m<n$).
Proof.
Recall that the the regularity of a sequence of homogeneous polynomials is equivalent to the semi-regularity if the number of polynomials is less than or equal to the number of variables. For the sequence $(f_{1},\ldots,f_{m},\ell)$, the number of polynomials and that of variables are both $n$, whence it is regular. Therefore, $(f_{1},\ldots,f_{m})$ is also regular, and
${\rm HS}_{R/I}(z)={\rm HS}_{R/\langle I,\ell\rangle}(z)\cdot\frac{1}{1-z}={\rm HS}_{R/\langle I,\ell\rangle}(z)\cdot(1+z+z^{2}+\cdots)$
@@ 3.1 Proofs of Theorem 1 (1) and (2)
${\rm HS}_{R/\langle I,\ell\rangle}(z)=\frac{\prod_{j=1}^{n-1}(1-z^{d_{j}})\cdot(1-z)}{(1-z)^{n}}=\frac{\prod_{j=1}^{n-1}(1-z^{d_{j}})}{(1-z)^{n-1}}=\prod_{j=1}^{n-1}(1+z+\cdots+z^{d_{j}-1}),$
@@
which is a polynomial. Here, we have $\mathrm{HF}_{R/I}(d)=\sum_{i=0}^{d}\mathrm{HF}_{R/\langle I,\ell\rangle}(i)$ for any $d$. Since $\mathrm{HF}_{R/\langle I,\ell\rangle}(d)=0$ (resp. $\mathrm{HF}_{R/\langle I,\ell\rangle}(d)\neq 0$) for $d\geq d_{\rm reg}(\langle I,\ell\rangle_{R})$ (resp. $d\leq d_{\rm reg}(\langle I,\ell\rangle_{R})-1$), we have $\widetilde{d}_{\rm reg}(I)=d_{\rm reg}(\langle I,\ell\rangle_{R})-1$, as desired. ∎
Proof of Theorem 1 (2):

### v2

@@ 2.3 Genericness and Fröberg’s conjecture
Since every Hilbert series is not changed by extending the field of coefficients, the Hilbert series of the quotient by an ideal generated by a generic sequence depends only on $\mathrm{char}(K)$, $n$, $m$, and $(d_{1},\ldots,d_{m})$, see e.g., [16, Lemma and Definition 4]. Thus, in discussing the Hilbert series, we may consider the ideal generated by $\bm{G}$ over $K_{0}(\mathcal{C})[X]$ instead of that over $K(\mathcal{C})[X]$. Because, the reduced Gröbner basis of $\langle\bm{G}\rangle_{R_{L}}$ coincides with that of $\langle\bm{G}\rangle_{K_{0}(\mathcal{C})[X]}$, since the well-known Buchberger’s algorithm can be performed over a field, over which the generating set is defined. Hence the minimal generating set of ${\rm in}(\langle\bm{G}\rangle_{R_{L}})$ coincides with that of ${\rm in}(\langle\bm{G}\rangle_{K_{0}(\mathcal{C})[X]})$. This also implies that, for computing the reduced Gröbner basis of $\langle\bm{G}\rangle_{R_{L}}$, we may assume that all elements of $\mathcal{C}$ are algebraically independent over $K$ and compute it over $L=K(\mathcal{C})$.
If $K$ is an infinite field, Fröberg-Löfwall proved in [26, Theorem 1] that there exists a non-empty Zariski-open subset $U_{\rm min}\subset V_{n,m,(d_{1},\ldots,d_{m})}$ on which the quotients by the corresponding ideals have the same Hilbert series, which is minimal in the coefficient-wise sense. Note that, even if $K$ contains the set $\mathcal{C}$ of $N$ algebraically independent elements, $\bm{F}\in U_{\rm min}$ not necessarily implies the genericness of $\bm{F}$.
On the other hand, the Hilbert series of $R/\langle\bm{F}\rangle$ for $\bm{F}\in U_{\rm min}$ is equal to that of $R_{L}/\langle\bm{G}\rangle$ for a generic sequence $\bm{G}$ over $L$ as in Lemma 2.3.2. The existence of the non-empty Zariski open set $U_{\rm min}$ can be easily verified by considering the stability of the reduced Gröbner basis in the context of so-called a comprehensive Gröbner system for a parametric ideal. A result by Nabeshima in [42] on “generic Gröbner basis of a parametric ideal” can be translated directly to our problem where the generic sequence $\bm{G}$ is given from $K[\mathcal{C}][x_{1},\ldots,x_{n}]$ and $\mathcal{C}$ is the set of coefficient parameters. Here we assume that all elements of $\mathcal{C}$ are algebraically independent over $K$ and hence they are treated as independent indeterminates. (See Remark 2.3.3.) In the below, we show a useful result for our purpose. We let $\prec_{1}$ be a block ordering with $X:=\{x_{1},\ldots,x_{n}\}\succ\succ\mathcal{C}$ such that its restriction on $X$ coincides with the given ordering $\prec$. Let $\bm{H}$ be a finite set of $K[\mathcal{C}][X]$ and ${\rm GB}(\bm{H})$ the reduced Gröbner basis of the ideal $\langle\bm{H}\rangle_{L[x]}$ with respect to the restriction $\prec$ of $\prec_{1}$ on the monomials in $X$, where $L=K(\mathcal{C})$. Moreover, let $\widetilde{\rm GB}(\bm{H}):=\{{\rm dlcm}(h)h:h\in{\rm GB}(\bm{H})\}$ and $\ell(\bm{H})=\prod_{h^{\prime}\in\widetilde{\rm GB}(\bm{H})}{\rm LC}_{\prec}(h^{\prime})$ in $K[\mathcal{C}][X]$, where ${\rm dlcm}(h)$ denotes the least common multiple of all denominators of coefficients in $K(\mathcal{C})$ of $h$ and each $h^{\prime}\in\widetilde{\rm GB}(\bm{H})$ is considered as an element in $K[\mathcal{C}][X]$, that is, a polynomial in $X$ over $K[\mathcal{C}]$. For $\bm{a}\in\mathbb{A}^{N}(K)$, we denote by $\pi_{\bm{a}}$ the substitution homomorphism from $K[\mathcal{C}][X]$ to $R=K[X]$, that is, $\pi_{\bm{a}}(h(\mathcal{C},X))=h(\bm{a},X)$ for $h\in K[\mathcal{C},X]$. For $h\in L[X]$, when the denominator of each coefficient of $h$ does not vanish at $\bm{a}$, the image $\pi_{\bm{a}}(h)$ is defined naturally.
Theorem 2.3.4 ([42, Theorem 3]).
With notation as above, consider $\bm{H}$ and $\widetilde{\rm GB}(\bm{H})$ to be subsets of $K[\mathcal{C}\cup X]$, and let $S(\bm{H})$ be the reduced Gröbner basis of the ideal quotient $\langle\bm{H}\rangle_{K[\mathcal{C}\cup X]}:\langle\widetilde{\rm GB}(\bm{H})\rangle_{K[\mathcal{C}\cup X]}$ with respect to $\prec_{1}$. Then, we have $S(\bm{H})\cap K[\mathcal{C}]\not=\emptyset$, and for any $\bm{a}$ in $\mathbb{A}^{N}(\overline{K})\smallsetminus(V_{\overline{K}}(S(\bm{H})\cap K[\mathcal{C}])\cup V_{\overline{K}}(\ell(\bm{H})))$, the set $\pi_{\bm{a}}({\rm GB}(\bm{H}))$ is the reduced Gröbner basis of $\langle\pi_{\bm{a}}(\bm{H})\rangle_{K[X]}$ with respect to $\prec$.
Now we return to our case where $\bm{H}=\bm{G}$. Let $\mathcal{O}=\mathbb{A}^{N}(K)\smallsetminus(V_{K}(S(\bm{G})\cap K[\mathcal{C}])\cup V_{K}(\ell(\bm{G})))$. Then $\mathcal{O}$ is Zariski-open in $\mathbb{A}^{N}(K)$. Moreover, $\mathcal{O}$ is not empty, since $\langle S(\bm{G})\cap K[C]\rangle\cap\langle\ell(\bm{G})\rangle$ is a non-zero ideal and since $K$ is an infinite field. It is obvious that for any $\bm{a}\in\mathcal{O}$, $\pi_{\bm{a}}({\rm GB}(\bm{G}))$ is the reduced Gröbner basis of $\langle\pi_{\bm{a}}(\bm{G})\rangle_{R}$ with respect to $\prec$. Moreover, for such $\bm{a}$, since ${\rm LC}_{\prec}(\pi_{\bm{a}}(h))=1$ for any $h\in{\rm GB}(\bm{G})$, we have ${\rm LM}_{\prec}(\pi_{\bm{a}}({\rm GB}(\bm{G})))={\rm LM}_{\prec}({\rm GB}(\bm{G}))$.
Corollary 2.3.5.
Under the assumption in Theorem 2.3.4, there is a non-empty Zariski open set $\mathcal{O}$ in $K^{N}$ such that, for any $\bm{a}$ in $\mathcal{O}$, the set of leading monomials of the reduced Gröbner basis of $\langle\pi_{\bm{a}}(\bm{G})\rangle_{R}$ coincides with that of $\langle\bm{G}\rangle_{R_{L}}$.
For each ideal $I$ of $K[X]$, the Hilbert series ${\rm HS}_{R/I}(z)$ is determined by its initial ideal and so by the leading monomials of its (reduced) Gröbner basis. Thus, Corollary 2.3.5 implies that ${\rm HS}_{R/\langle\pi_{\bm{a}}(\bm{G})\rangle}(z)$ coincides with ${\rm HS}_{R_{L}/\langle\bm{G}\rangle}(z)$ for any $\bm{a}\in\mathcal{O}$. Considering the corresponding set $U$ of $\mathcal{O}$ in $V_{n,m,(d_{1},\ldots,d_{m})}$, we have the following proposition and a conjecture. Here, we remark that the following proposition may be a well-known fact. However, the authors could not find any reference for its proof, and therefore they would write a proof here for the readers’ convenience; the idea of the proof will be used to obtain an analogous result in Section 4 below.
Proposition 2.3.6.
Assume that $K$ is an infinite field. Then, for given $n$, $m$, and $(d_{1},\ldots,d_{m})$, the following conditions (1), (2), (3), and (4) are equivalent to each other:
  1. (1)
    There exists an $\bm{F}=(f_{1},\ldots,f_{m})\in V_{n,m,(d_{1},\ldots,d_{m})}$ satisfying (2.2.2), i.e., $\bm{F}$ is cryptographic semi-regular (in other words, the ideal $\langle\bm{F}\rangle$ is sufficiently generic).
  2. (2)
    For the non-empty Zariski open set $U_{\rm min}$ defined in the paragraph just after Remark 2.3.3, any $\bm{F}=(f_{1},\ldots,f_{m})\in U_{\rm min}$ satisfies (2.2.2).
  3. (3)
    The property that $\bm{F}=(f_{1},\ldots,f_{m})\in V_{n,m,(d_{1},\ldots,d_{m})}$ satisfies (2.2.2) is generic (in terms of Definition 2.3.1).
  4. (4)
    For $\mathcal{C}$, $L$, $R_{L}$, and $\bm{G}=(g_{1},\ldots,g_{m})$ as in Lemma 2.3.2, we have
    $\mathrm{HS}_{R_{L}/\langle g_{1},\ldots,g_{m}\rangle}(z)=\left[\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\right].$ (2.3.2)
    Namely, the generic sequence $\bm{G}$ is cryptographic semi-regular (in other words, the ideal $\langle\bm{G}\rangle_{R_{L}}$ is sufficiently generic).
Proof.
The implication (2) $\Rightarrow$ (1) is obvious.
To see (1) $\Rightarrow$ (4), it follows from Lemma 2.2.6 and Lemma 2.3.2 that
$\mathrm{HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}(z)\geq\mathrm{HS}_{R_{L}/\langle g_{1},\ldots,g_{m}\rangle}(z)\geq^{\rm lex}\left[\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\right].$ (2.3.3)
By our assumption that $\mathrm{HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}(z)$ satisfies (2.2.2), we have (2.3.2).
The implication (4) $\Rightarrow$ (3) follows from Corollary 2.3.5.
The implication (3) $\Rightarrow$ (2) follows from the irreducibility of $V_{n,m,(d_{1},\ldots,d_{m})}$. ∎
@@
Assume that $K$ is an infinite field. For any $n$, $m$, and $(d_{1},\ldots,d_{m})$, the equivalent conditions (1), (2), (3), and (4) in Proposition 2.3.6 hold.
See Pardue’s paper [43] for some conjectures equivalent to Fröberg’s one (Conjecture 2.3.7). He also proved that Moreno-Socías conjecture [41] is stronger than Conjecture 2.3.7, see [43, Theorem 2] for a proof.
3 Proofs of Theorems 1.1 and 1.2
As in the previous sections, let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$ (in any characteristic). Let $I$ be a proper homogeneous ideal of $R$, and let $F=\{f_{1},\ldots,f_{m}\}$ be a homogeneous subset of $R$ generating $I$. For each $1\leq j\leq m$, put $d_{j}:=\deg(f_{j})$ (total degree).
3.1 Proof of Theorem 1.1
We first prove Theorem 1.1 (1), which can be easily proved by Lemma 2.2.11 together with the following easy lemma in the case where $m=n-1$:
@@
With notation as above, if $m=n-1$ and if the sequence $(f_{1},\ldots,f_{m},x_{n})$ is cryptographic semi-regular (equivalently, it is regular since $m+1=n$), then we have $\widetilde{d}_{\rm reg}(I)=d_{\rm reg}(I+\langle x_{n}\rangle_{R})-1$, and hence
$\mathrm{max.GB.deg}_{\prec}(I)\leq\mathrm{max}\{d_{\rm reg}(I+\langle x_{n}\rangle),\widetilde{d}_{\rm reg}(I)\}=d_{\rm reg}(I+\langle x_{n}\rangle)=\sum_{j=1}^{n-1}(d_{j}-1)+1$
for the graded reverse lexicographical ordering $\prec$ with $x_{n}\prec\cdots\prec x_{1}$. Moreover, $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular (equivalently, it is regular since $m<n$).
@@
Since $(f_{1},\ldots,f_{m},x_{n})$ is regular, $(f_{1},\ldots,f_{m})$ is also regular, so that
${\rm HS}_{R/I}(z)={\rm HS}_{R/(I+\langle x_{n}\rangle)}(z)\cdot\frac{1}{1-z}={\rm HS}_{R/(I+\langle x_{n}\rangle)}(z)\cdot(1+z+z^{2}+\cdots)$
@@
${\rm HS}_{R/(I+\langle x_{n}\rangle)}(z)=\frac{\prod_{j=1}^{n-1}(1-z^{d_{j}})\cdot(1-z)}{(1-z)^{n}}=\frac{\prod_{j=1}^{n-1}(1-z^{d_{j}})}{(1-z)^{n-1}}=\prod_{j=1}^{n-1}\left(\sum_{k=0}^{d_{j}-1}z^{k}\right),$
@@
which is a polynomial. It is clear that $\widetilde{d}_{\rm reg}(I)=d_{\rm reg}(I+\langle x_{n}\rangle_{R})-1$, as desired. ∎
Proof of Theorem 1.1 (1): If $m\geq n$, our assumption that $(f_{1},\ldots,f_{m},x_{n})$ is cryptographic semi-regular implies
${\color[rgb]{0,0,0}d_{\rm reg}(I+\langle x_{n}\rangle)=\deg\left(\left[\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n-1}}\right]\right)+1,}$